A sample​ mean, sample​ size, population standard​ deviation, and confidence level are provided. Use this information to complete parts​ (a) through​ (c) below.

x overbarx equals=25​,

n equals=38​,

sigma σ equals=4

confidence

level equals=95%

Click here to view page 1 of the standard normal distribution table.

LOADING...

Click here to view page 2 of the standard normal distribution table.

LOADING...

. Use the​ one-mean z-interval procedure to find a confidence interval for the mean of the population from which the sample was drawn.

The confidence interval is from

________

to

_______.
Answer #1

The z-value for 95% confidence interval is 1.96
Sample mean = 25
Standard deviation = 4
Standard error of mean = σ / √ n
Standard error of mean = 4 / √ 38 = 0.65

95 % confidence interval = 25-(0.65)*(1.96) and  25+(0.65)*(1.96)

=(23.73, 26.27)

The confidence interval is from 23.73 to 26.27